9. A company sells a container of mixed nuts that is 12% peanuts and another container that is
60% peanuts. How many cups of each mixture would be needed to make 12 cups that is 40%
peanuts?
4 cups of the 12% peanuts and 8 cups of the 60% peanuts
8 cups of the 12% peanuts and 4 cups of the 60% peanuts
5 cups of the 12% peanuts and 7 cups of the 60% peanuts
3 cups of the 12% peanuts and 9 cups of the 60% peanuts

Found 2 solutions by jorel1380, Boreal:
Answer by jorel1380(3719)

Let n be the amount of 60% peanuts. Then the amount of 12% peanuts is 12-n. So:
.6n+.12(12-n)=.4(12)
.48n+1.44=4.8
.48n=3.36
n=7 cups 60% peanuts
12-n=5 cups 12% peanuts. ☺☺☺☺

Answer by Boreal(15235)

x= cups of 12%
12-x= cups of 60%
.12x+.60(12-x)=.40(12)
.12x+7.2-.60x=4.8
-.48x=-2.4
x=5 cups of 12%
12-x=7 cups of 60%
expect the number of cups of 60% to be more, because you end up a little closer to 60% than to 12%.